OpenStudy (anonymous):

x + 5y = -2 2x + y = 5 The point of intersection of the lines has a y-coordinate of _____.

OpenStudy (mehek.15):

would you like to solve by substitution or elimination?

OpenStudy (anonymous):

Elimination

OpenStudy (anonymous):

The options are -1 .. 1 .. 3

OpenStudy (mehek.15):

oke multiply first equation by 2

OpenStudy (anonymous):

ok one sec

OpenStudy (anonymous):

pls don't leave lol

OpenStudy (mehek.15):

I wont :P

OpenStudy (anonymous):

K. I got 2x + 10y = -4

OpenStudy (mehek.15):

subtract 2x+10y = -4 - 2x + y = 5 -----------------

OpenStudy (anonymous):

How do I know when to subtract and when to add the equations?

OpenStudy (anonymous):

or*

OpenStudy (mehek.15):

2x - 2x = ? 10y - y = ? -4 - 5 = ?

OpenStudy (anonymous):

Yeah... like how u said, subtract. If I added, my answer would be incorrect?

OpenStudy (mehek.15):

ye subtracting = eliminating x because 2x - 2x = 0

OpenStudy (anonymous):

Oh. Because my lesson says that you can add as long as you do it on both sides of the equation

OpenStudy (anonymous):

Lol. This is confusing. But anyway, my answer was 1 :)

OpenStudy (mehek.15):

if u add in this case, none will be eliminated

OpenStudy (anonymous):

Ohh. Yeah, you're right.

OpenStudy (anonymous):

what grade are you in btw? I'm in 10th. And thanks for the help.

OpenStudy (mehek.15):

9th

OpenStudy (mehek.15):

and the answer is not 1 :P

OpenStudy (anonymous):

Oh wow. Lol, I'm just learning this topic as we speak

OpenStudy (anonymous):

but -4 +5 is 1

OpenStudy (anonymous):

2x + 10y = -4 2x + y = 5 0x = 1

OpenStudy (mehek.15):

2x - 2x = 0 10y - y = 9y -4 - 5 = -9 9y = -9 y = ?

OpenStudy (anonymous):

ohhhh. -1 lol

OpenStudy (mehek.15):

ye :P

OpenStudy (anonymous):

Great! Hopefully I understand the rest of my questions. Thankyouuuu

OpenStudy (mehek.15):

yw :)